# LLMs Homogenize Values in Constructive Arguments on Value-Laden Topics

### v1

@@ 3.3. Phase 3: Perceived Value Alignment
Experiment Design. After collecting 218 pairs of human-written and LLM-written comments from both Phase 1 and Phase 2, we conducted a forced-choice experiment to examine what types of comments people perceived as more aligned with their values. We recruited 92 Indian and 92 American crowd workers from Prolific, who had not participated in both Phase 1 and Phase 2. We initially asked participants about their stance on same-sex marriage and Islam. Then they were randomly assigned to review comments either on homophobic or Islamophobic threads. Each participant reviewed four randomly selected pairs of human-written comments and their corresponding LLM-rewritten versions. We ensured that each participant reviewed pairs of comments in which the human-written version was authored by participants from the same country (in Phase 1 and 2) and shared the same stance as the reviewing participant. This is because, a participant supportive of same-sex marriage or Islam may not find comments written against these topics aligned with their values and vice versa.
We randomized the order of comments within each pair and asked participants to select which comment aligned more with their values. We did not reveal to participants about how these comments were written. We also asked them open-ended questions about how their chosen comments aligned with their values. In total, we received 1122 human evaluations from 180 participants after discarding responses from four participants, who failed the attention check. Participants were compensated with USD 1.00 for completing the task.
Data Analysis. We conducted multiple Chi-square tests with Bonferroni correction to examine whether participants’ preferences for human-written vs. LLM-rewritten comments varied depending on their stance on same-sex marriage and Islam. We also reviewed the open-ended responses they provided for how their chosen comments aligned with their values.
@@ 4. Findings
We first describe results from Phase 1 to answer what values Indians and Americans expressed when writing constructive comments on homophobic and Islamophobic threads (4.1). Then from Phase 2, we present our analyses of what values LLM emphasized when rewriting comments constructively on value-laden issues (4.2). Finally, we outline what types of comment people found more aligned with their values (4.3).
@@
4.1. RQ1: What Values Do People Emphasize When Writing Constructively on Value-Laden Topics?
@@
Our annotation of 116 human-written constructive comments resulted in 681 segments containing Schwartz’s basic human values. On average, we identified five underlying values (SD=2.17) in each comment.
@@
Figure 5. Distribution of Schwartz’s human values in (A) all comments written by Indian and American participants, (B) comments written on homophobic thread by participants who either support or oppose same-sex marriage. Statistically significant differences are reported at $p<0.0001$ (***), $p<0.001$ (**), and $p<0.01$ (*)[adjusted P-values after Bonferroni correction].
[…]
@@
A chi-square test with Yates’ continuity correction revealed significant cross-cultural differences in the distribution of Schwartz’s basic human values between constructive comments written by Indian and American participants; though the effect size was small $(\chi^{2}(7,N=675)=24.32,p<0.001,\phi=0.19$). Post hoc analysis with Bonferroni correction showed that Indian participants (13%) expressed significantly more Traditional values than the American participants (5%) when writing constructively on the same threads (see Figure 5A). For example, in response to the homophobic threads, some Indian participants invoked concerns about how same-sex marriages might disrupt the socio-cultural, ancestral, and religious traditions. On Islamophobic threads, these comments referred to different traditions practiced by Muslims. This pattern is consistent with prior research showing that Indian society places greater emphasis on traditional values than American society (Sundberg et al., 1970).
@@
Cross-Cultural Values Across Topic. When we analyzed the underlying values across topics, we observed that the distribution of values differed significantly between constructive comments written by Indian and American participants; but only for the homophobic thread with a small effect size $(\chi^{2}(7,N=355)=18.52,p<0.001,\phi=0.23$). Post hoc analysis with Bonferroni correction showed that Indians (19.7%) expressed significantly more Traditional values than the Americans (6.4%) when writing constructive comments on the homophobic thread. We did not notice any significant cross-cultural difference for the Islamophobic thread. This indicates that the observed cross-cultural differences in values might be linked to participants’ stances on same-sex marriage, which we examine next.
@@
Cross-Cultural Values Across Stance. Indian and American participants’ stances on same-sex marriage differed significantly; with a medium effect size $(\chi^{2}(2,N=58)=10.4,p<0.01,\phi=0.42$). Post hoc analysis with Bonferroni correction showed that Americans supported same-sex marriage (72%) significantly more than the Indians (31%) (see Table 1). In contrast, no significant cultural differences emerged in participants’ stances toward Islam, which may explain the absence of cross-cultural differences in the underlying values expressed in comments on the Islamophobic thread.
@@
Against: 28%
Neutral: 41%
Support: 31%
Against: 7%
Neutral: 21%
Support: 72%
Against: 24%
Neutral: 28%
Support: 48%
Against: 38%
@@
Support: 38%
To further examine how culture and stance shaped the underlying values in constructive comments on the homophobic thread, we fit a multinomial logistic regression model. The analysis revealed only a significant main effect of stance on the value expressions $(\chi^{2}(7,N=355)=34.57,p<0.0001,\phi=0.37$). Post hoc analysis with Bonferroni correction also showed that participants who opposed same-sex marriage expressed significantly more (31.3%) Traditional values in their comments than those who supported (7.2%) such marriage (see Figure 5B). Table 2 gives an example of how opponents of same-sex marriage framed their arguments from a Traditional point of view.
Table 2. Examples of constructive comments written on homophobic thread by an Indian participant opposing same-sex marriage and an American participant supporting such marriage, with highlighted segments illustrating Traditional values.
@@
Marriage is a sacred practice. It is not
that simple like between two body it
is much more beyond that it is the
unification of two souls as they have
to share everything their happiness
their sorrow all of these. It is only
possible between a man and woman.
In today’s generation they think
marriage is like a sale whenever they
wanted they marry and whenever they
want they got divorce. Gay is being
normalised but it is not good for the
society to function. God has created
both men and women so they can
share everything between each other.
So in my opinion marriage should be
only between men and women.
@@
their lives however they choose as long
as they are not hurting anybody else.
Don’t we want everyone to be happy?
That involves allowing people to marry
whoever they want, assuming it is legal
in other aspects as well. People who
oppose same sex marriage don’t have
anything better to do with their lives
but judge others and be mad all the time.
Overall, these findings indicate that people’s stances on same-sex marriage, rather than culture, influenced the values they emphasized when writing constructively on homophobic threads.
@@ 4.2. RQ2: What Values Do LLMs Prioritize When Writing Constructively on Value-Laden Topics?
Are LLM-Rewrites Constructive?. We collected 218 pairs of human-written and LLM-rewritten comments both from Phase 1 and Phase 2. Our analyses revealed that these LLM-rewritten versions had significantly more linguistic features of constructiveness than original human-written comments (see Table 8 in Appendix). For example, LLM-rewritten comments were longer, more readable, more polite, more argumentative, i.e., contained more discourse connectives (logical reasoning in argument), reasoning verbs (conclusion in argument), modals (proposition in argument), and root clauses (independent clause) than the original human-written comments. This aligns with prior finding that LLMs can rewrite comments more constructively than humans (Shahid et al., 2025).
Do LLM-Rewrites Have Same/ Different Values?. Our annotation of 218 pairs of human-written and LLM-rewritten constructive comments led to 2218 comment segments containing Schwartz’s basic human values. On average, we identified five values (SD=2.18) within each comment. Figure 6 shows the distribution of values present across human-written and LLM-rewritten constructive comments.
Table 3. Example of a constructive comment written by an American participant on an Islamophobic thread, alongside the corresponding LLM-rewritten version, with highlighted segments illustrating the values of Security, Conformity, and Benevolence.
@@
There is a lot in the media that portrays
Muslims as barbaric and aggressive
so it is an understandable viewpoint
whether it is true or not. There
continues to be media coverage
showing hostage situations and brutal
killings that would easily paint that
picture in people’s minds and sway
their viewpoint about Muslims in
general when that behavior is most
likely limited to specific terrorist groups.
The media often highlights violent incidents
involving specific terrorist groups, which
may inadvertently shape perceptions about
Muslims as a whole. It’s important to
differentiate between these groups and the
vast majority of Muslims who do not engage
in such actions. Encouraging media outlets to
provide a balanced perspective, highlighting
the diverse contributions and peaceful practices
of Muslim communities, can help counteract
these negative portrayals. Open dialogues and
educational initiatives can also foster better
understanding and reduce stereotypes.
@@
We observed significant differences in the distribution of values between human-written and LLM-rewritten constructive comments, with a small effect size $(\chi^{2}(9,N=1329)=112.53,p<0.000001,\phi=0.23$). Post hoc tests with Bonferroni correction showed that LLMs downplayed values associated with Security (concern for social order and stability; human: 15.62%, LLM: 9.63%) and Conformity (adherence to rules; human: 5.95%, LLM: 2.29%), while emphasizing Benevolence (concern for others’ welfare; human: 9.58%, LLM: 18.78%) and Universalism (concern for equality and social justice; human: 33.27%, LLM: 44.57%) when rewriting people’s comments. Table 3 gives example of how LLM enhances Benevolence in its rewrite while downplaying Security concern and rule Conformity present in the original comment.
Topic Cases Statistics Value Changes
Culture India $\chi^{2}(9,N=1128)=67.26,p<0.000001,\phi=0.24$
$\downarrow$ Conformity (human: 5.9%, LLM: 1.6%)
$\uparrow$ Universalism (human: 33%. LLM: 46.6%)
US $\chi^{2}(9,N=1089)=54.2,p<0.000001,\phi=0.22$
$\downarrow$ Self-direction (human: 26.9%, LLM: 17.1%)
$\uparrow$ Benevolence (human: 8.5%. LLM: 19.5%)
Topic Homophobia $\chi^{2}(9,N=1126)=71.12,p<0.000001,\phi=0.25$ $\uparrow$ Benevolence (human: 5.4%. LLM: 16.2%)
Islamophobia $\chi^{2}(9,N=1089)=49.44,p<0.000001,\phi=0.21$
$\downarrow$ Security (human: 21.6%, LLM: 12.3%)
$\uparrow$ Universalism (human: 32.7%. LLM: 44.4%)
Stance Against same-sex marriage $\chi^{2}(6,N=263)=43.34,p<0.000001,\phi=0.41$
$\downarrow$ Tradition (human: 27.4%, LLM: 8%)
$\uparrow$ Universalism (human: 19.5%, LLM: 47.3%)
Supports same-sex marriage - -
Against Islam $\chi^{2}(7,N=371)=23.12,p<0.001,\phi=0.25$ $\uparrow$ Universalism (human: 26.5%, LLM: 44.4%)
Supports Islam - -
@@
Subsequent Chi-square tests with Bonferroni correction showed that LLM consistently shifted underlying values when rewriting people’s comments constructively across different cultures, topics, and stances; though effect sizes varied from small to medium (see Table 4). For instance, although Indians are known to exhibit greater Conformity to rules and social order (Sundberg et al., 1970; Anderson, 2012), LLM downplayed this value when rewriting their comments. In contrast, while Americans are known to prioritize Self-direction (i.e., freedom of thought and action) (Ryckman and Houston, 2003), LLM reduced emphasis on this value. Similarly, in response to Islamophobic threads, many participants referred to Security concern around terrorism and extremism, which LLM deprioritized during rewriting. On the other hand, as observed in Phase 1 (see Figure 5B), opponents of same-sex marriage frequently appealed to Traditional values, which the LLM minimized in its rewriting.
@@
According to Schwartz’s theory (Schwartz, 1994), the values that LLM consistently downplayed—Conformity, Security, and Tradition—fall under the higher-order value of Conservation, which emphasizes stability and the preservation of the existing social order. In contrast, Benevolence and Universalism—the values that LLM prioritized—are part of the higher-order value of Self-Transcendence, which prioritizes the welfare of others and the society.
@@
Notably, participants who wrote comments supporting same-sex marriage and Islam already emphasized Benevolence or Universalism by appealing to equality and rejecting dehumanization of these groups. This might explain why LLM did not shift underlying values significantly when rewriting comments from participants with prosocial viewpoints (see Table 4). Moreover, our manual inspection revealed that for comments authored by opponents of same-sex marriage and Islam, changes in values in LLM-rewritten versions frequently (89% of cases) led to corresponding changes in stance—moving them toward more neutral or supportive positions. For instance, the example in Table 5 shows that although the participant expressed their views against same-sex marriage, LLM totally altered their arguments when rewriting constructively. In contrast, comments from those supportive did not undergo such changes.
@@
having gay marriages or legalizing it
@@
biologically there are only two
genders male and female. I don’t think
gay marriages should be legalized in
India as marriages are meant for
having a child which is not possible
with gay marriages and it will not
only affect the heritages but also
the culture of India.
It’s important to understand that marriage
is not exclusively meant for procreation.
Many heterosexual couples also choose
not to have children or can’t have them
due to biological reasons. However, they
are still allowed to marry. Legalizing
same-sex marriages wouldn’t necessarily
impact India’s culture or heritage, as it
merely extends the legal rights and
protections of marriage to all citizens.
Moreover, gender diversity is biologically
proven beyond just male and female
categories, adding another dimension to
our understanding of human nature.

### v2

@@ 3.3. Phase 3: Perceived Value Alignment
Experiment Design. After collecting human-written and LLM-written comments from both Phase 1 and Phase 2, we conducted a forced-choice experiment to examine what types of comments people perceived as more aligned with their values. We recruited 92 Indian and 92 American crowd workers from Prolific, who had not participated in both Phase 1 and Phase 2. We initially asked participants about their stance on same-sex marriage and Islam. Then they were randomly assigned to review comments either on homophobic or Islamophobic threads. Each participant reviewed four randomly selected pairs of human-written comments and their corresponding LLM-rewritten versions. Since LLMs were prompted to rewrite constructively, we selected comments from the Phase 1 constructive condition (N=116) where participants were explicitly instructed to write constructively—to ensure that both comments in a pair were generated under comparable instructions. We controlled for country and stance, ensuring that each participant in Phase 3 reviewed comment pairs that matched their stance and were written by participants from their country (in Phase 1). This was necessary because participants who support same-sex marriage or Islam may not view comments written against these topics as aligned with their values, and vice versa.
We randomized the order of comments within each pair and asked participants to select which comment aligned more with their values. We did not reveal to participants how these comments were written. We also asked them open-ended questions about how their chosen comments aligned with their values. In total, we received 1122 human evaluations from 180 participants after discarding responses from four participants, who failed the attention check. Participants received USD 1.00 for completing the task.
Data Analysis. We conducted multiple CHI-square tests with Bonferroni correction to examine whether participants’ preferences for human-written vs. LLM-rewritten comments varied depending on their stance on same-sex marriage and Islam. We also conducted reflexive thematic analysis (Braun and Clarke, 2006) on participants’ open-ended responses describing reasons for their choice of constructive comments. We performed iterative open coding and revised the codes through discussion, resulting in 26 unique codes grouped into three higher-level themes (e.g., personal value, linguistic property, and extreme opposition). See Table 11 in Appendix for more details.
@@ 4. Findings
We first describe results from Phase 1 to answer what values Indians and Americans expressed when writing comments on homophobic and Islamophobic threads (4.1). Then from Phase 2, we present our analyses of what values LLM emphasized when rewriting comments constructively on value-laden issues (4.2). Finally from Phase 3, we outline what types of comment people found more aligned with their values (4.3).
@@
4.1. RQ1: What Values Do People Emphasize When Writing Comments on Value-Laden Topics?
@@
In Phase 1, we collected a total of 234 comments from Indian and American participants. Our annotation of these comments yielded 1,205 segments coded with Schwartz’s basic human values. On average, we identified five values (SD=1.95) in each comment.
@@
Figure 6. Distribution of Schwartz’s human values in (A) all comments written by Indian and American participants, (B) comments written on homophobic thread by participants who either support or oppose same-sex marriage. Statistically significant differences are reported at $p<0.00001$ (****), $p<0.0001$ (***), $p<0.001$ (**), and $p<0.01$ (*)[adjusted P-values after Bonferroni correction].
(A) Comments written by Indian participants express significantly more Traditional values than that of the Americans. (B) Comments written on homophobic thread by participants opposing same-sex marriage express significantly more Traditional values than those who support same-sex marriage. Grouped bar charts illustrating the distribution of Schwartz’s human values in: (A) all comments written by Indian and American participants, (B) comments written on homophobic thread by participants who either support or oppose same-sex marriage. Statistically significant differences exist only for Traditional values in both cases.
@@
Values Across Cultures. For our analyses, we considered participant’s country as a proxy for their culture. A CHI-square test with Yates’ continuity correction revealed significant cross-cultural differences in the distribution of Schwartz’s basic human values between comments written by Indian and American participants; though the effect size was small $(\chi^{2}(7,N=1195)=39.15,p<0.00001,\phi=0.18$). Post hoc analysis with Bonferroni correction showed that Indian participants (14%) expressed significantly more Traditional values than the American participants (6%) (see Figure 6A). This pattern is consistent with prior research showing that Indian society places greater emphasis on traditional values than the American society (Sundberg et al., 1970).
@@
Cross-Cultural Values Across Topic. When we examined the values within specific topics, we noticed that the cross-cultural differences only emerged for the homophobic thread, but the effect size was small $(\chi^{2}(6,N=606)=25.15,p<0.001,\phi=0.20$). Posthoc analysis with Bonferroni correction showed that Indian participants (22.7%) expressed significantly more Traditional values than the American participants (9.1%) when commenting on the homophobic thread. Many Indian participants invoked concerns that same-sex marriage would disrupt socio-cultural, ancestral, and religious traditions in India. In contrast, we did not notice any significant cross-cultural difference for the Islamophobic thread. This indicates that the observed cross-cultural differences in values might be linked to participants’ stances on same-sex marriage.
@@
Cross-Cultural Values Across Stance. Indian and American participants’ stances on same-sex marriage differed significantly; with a medium effect size $(\chi^{2}(2,N=117)=25.96,p<0.00001,\phi=0.47$). Post hoc analysis with Bonferroni correction showed that Americans supported same-sex marriage (72%) significantly more than the Indians (27%), whereas Indian participants opposed same-sex marriage (44%) significantly more than their American counterparts (10%) (see Table 2). In contrast, participants did not differ in their views on Islam, which may help explain why no cross-cultural differences emerged in the values expressed in comments on the Islamophobic thread.
@@
Against: 44%
Neutral: 29%
Support: 27%
Against: 10%
Neutral: 17%
Support: 73%
Against: 39%
Neutral: 20%
Support: 41%
Against: 36%
@@
Support: 40%
Table 3. Examples of comments written on homophobic thread by an Indian participant opposing same-sex marriage and an American participant supporting such marriage, with highlighted segments illustrating Traditional values.
@@
Marriage is a sacred practice. It is not that simple like between two
body it is much more beyond that. it is the unification of two souls
as they have to share everything their happiness their sorrow all of
these. It is only possible between a man and woman. In today’s
generation they think marriage is like a sale whenever they
wanted they marry and whenever they want they got divorce. Gay
is being normalised but it is not good for the society to function.
God has created both men and women so they can share everything
between each other. So in my opinion marriage should be only
between men and women.
@@
their lives however they choose as long as
they are not hurting anybody else. Don’t we
want everyone to be happy? That involves
allowing people to marry whoever they
want, assuming it is legal in other aspects
as well. People who oppose same sex
marriage don’t have anything better to do
with their lives but judge others and be mad
all the time.
To further examine how participant’s culture and stance shaped their expression of values in comments on the homophobic thread, we fit a multinomial logistic regression model. The analysis revealed only a significant main effect of stance on the value expressions $(\chi^{2}(7,N=462)=58.1,p<0.000001,\phi=0.35$). Post hoc analysis with Bonferroni correction also showed that participants who opposed same-sex marriage expressed significantly more (31.9%) Traditional values in their comments than those who supported (8.9%) such marriage (see Figure 6B). Table 3 gives an example of how opponents of same-sex marriage framed their arguments from a Traditional point of view.
Values Across Different Demographics. Table 4 lists the demographic details of our participants. Indian and American participants differed significantly in terms of their age ($\chi^{2}(2,N=117)=11.07,p<0.01,\phi=0.31$) and religion ($\chi^{2}(11,N=117)=98.39,p<0.000001,\phi=0.92$). Majority of the Indian participants (76%) were young, whereas most of the American participants were either young or middle aged. On the other hand, majority of the Indian participants (76%) identified as Hindu while those in the US primarily identified as either Christian (47%) or Atheist (40%). We did not notice any significant difference in the gender composition of our participants.
Table 4. Demographic details of Indian and American participants in Phase 1. We excluded the religious categories that had only one or two participants.
India United States
Age Gender Religion Age Gender Religion
Young (18-39): 76%
Middle (40-59): 24%
Old (60+): 0%
Female: 49%
Male: 50%
Hindu: 76%
Muslim: 10%
Christian: 3%
Atheist: 3%
Young (18-39): 48%
Middle (40-59): 47%
Old (60+): 5%
Female: 60%
Male: 40%
Hindu: 0%
Muslim: 0%
Christian: 47%
Atheist: 40%
When we analyzed the distribution of values, we did not observe any significant difference in the comments written by different age and gender groups. However, we observed significant but small effects of religion on the values participants expressed in their comments. For example, in Islamophobic thread, Muslim participants expressed significantly more Traditional values (15.4%) than other religious groups (2-4%) ($\chi^{2}(18,N=529)=49.38,p<0.0001,\phi=0.18$). They highlighted the peaceful teachings of Islam to counter Islamophobic claims. In contrast, Christian participants responded to Islamophobic claims by emphasizing Universalism (tolerance and acceptance of others) at a significantly higher rate (42%) than others (Muslims: 12%, Hindus: 27%, Atheist: 32%). They highlighted that it is inappropriate to generalize against Muslims, noting that other organized religions also have histories of violence.
On the other hand, Hindu participants expressed significantly more Traditional values when commenting on homophobic thread; but with a small effect size ($\chi^{2}(18,N=529)=41.38,p<0.01,\phi=0.16$). This pattern is consistent with our earlier finding that Indian participants expressed more Traditional values than American participants, as all Hindu participants in our sample were from India. However, when we fit a multinomial logistic regression model to examine how culture and religion may affect one’s value expression on these threads, we did not find any significant effect of religion.
Overall, these findings indicate that Indian and American participants’ divergent stances on same-sex marriage, rather than culture or demographic factors (age, gender, and religion) influenced the values they emphasized when commenting on homophobic thread.
@@ 4.2. RQ2: What Values Do LLMs Prioritize When Writing Constructively on Value-Laden Topics?
Are LLM-Rewrites Constructive?. We collected 336 pairs of human-written and LLM-rewritten comments both from Phase 1 and Phase 2. Our analyses revealed that these LLM-rewritten versions had significantly more linguistic features of constructiveness than original human-written comments (see Table 5). For example, LLM-rewritten comments were longer, more readable, more polite, more argumentative, i.e., contained more reasoning verbs (conclusion in argument), modals (proposition in argument), and root clauses (independent clause) than the original human-written comments. This aligns with prior finding that LLMs can rewrite comments more constructively than humans (Shahid et al., 2025).
Feature Statistics Human-written (Average) LLM-rewritten (Average)
Length W=7316.5, Z=-11.37, p<0.000001, r=0.44 66.1 83.1
Readability W=260.5, Z=-15.54, p<0.000001, r=0.61 7.9 15.4
Named entity W=13617, Z=6.21, p<0.000001, r=0.24 2.2 1.6
Discourse connective - 6.5 6.4
Stance adverbial W=10024, Z=5.43, p<0.000001, r=0.21 0.64 0.32
Reasoning verbs and modals W=2753, Z=-13.39, p<0.000001, r=0.53 4.7 7.7
Root clauses W=5953, Z=9.82, p<0.000001, r=0.38 3.7 4.7
Politeness W=14752, Z=-5.93, p<0.000001, r=0.23 13.17 15.17
Do LLM-Rewrites Have Same/ Different Values?. Our annotation of 336 pairs of human-written and LLM-rewritten constructive comments led to 3,351 segments containing Schwartz’s basic human values. On average, we identified five values (SD=2.29) within each comment. Figure 7 shows the distribution of values across human-written and LLM-rewritten constructive comments.
Table 6. Example of a comment written by an American participant on an Islamophobic thread, alongside the corresponding LLM-rewritten constructive version, with highlighted segments illustrating the values of Security and Benevolence.
@@
Islam has a long history of violence since it’s
inception. Islamic terrorism is a real thing that
is condoned by Muslims. Jihadists have killed
a great number of people. Don’t you think this
points to the fact that this religion is prone to
violence? It is just a fact that there is more
violence among Muslims than any other religion.
Throughout history, many religions, including Islam, have faced challenges with
individuals or groups engaging in violence. It’s important to separate extremist
actions from the beliefs of the broader community. While some groups have
committed violence in the name of religion, the vast majority of Muslims around
the world seek peace and promote understanding. Rather than attributing these
actions to Islam as a whole, it would be more productive to explore and address
the root causes of extremism and work together to foster peace and unity across
all communities.
@@
We observed significant differences in the distribution of values between human-written and LLM-rewritten constructive comments, with a small effect size $(\chi^{2}(9,N=3348)=201.73,p<0.000001,\phi=0.24$). Post hoc tests with Bonferroni correction showed that LLMs downplayed values associated with Security (concern for social order and stability; human: 18.21%, LLM: 10.64%), Conformity (adherence to rules; human: 6.33%, LLM: 2.35%), and Tradition (commitment to one’s culture or religion; human: 10.03%, LLM: 5.02%), while emphasizing Benevolence (concern for others’ welfare; human: 12.07%, LLM: 21.23%) and Universalism (concern for equality and social justice; human: 29.62%, LLM: 42.69%) when rewriting people’s comments. Table 6 gives an example of how the LLM shifts the framing to Benevolence while softening Security concern present in the original comment.
Factor Cases Statistics Value Changes
@@
$\downarrow$ Conformity (human: 6.9%, LLM: 1.7%)
$\downarrow$ Security (human: 20.7%, LLM: 9.9%)
$\downarrow$ Tradition (human: 13.5%, LLM: 6.3%)
[…]
$\uparrow$ Benevolence (human: 12.4%. LLM: 22.0%)
$\uparrow$ Universalism (human: 27.1%. LLM: 43.7%)
[…]
$\uparrow$ Universalism (human: 32.4%, LLM: 41.7%)
$\uparrow$ Benevolence (human: 11.8%. LLM: 20.5%)
[…]
$\downarrow$ Conformity (human: 8.0%, LLM: 2.9%)
$\downarrow$ Tradition (human: 15.1%, LLM: 6.6%)
[…]
$\uparrow$ Benevolence (human: 8.2%. LLM: 19.5%)
$\uparrow$ Universalism (human: 31.0%. LLM: 43.4%)
[…]
$\uparrow$ Benevolence (human: 16.0%. LLM: 23.1%)
$\uparrow$ Universalism (human: 28.4%. LLM: 42.1%)
[…]
$\downarrow$ Conformity (human: 12.4%, LLM: 3.1%)
$\downarrow$ Security (human: 14.4%, LLM: 3.1%)
$\downarrow$ Tradition (human: 29.9%, LLM: 9.4%)
[…]
$\uparrow$ Benevolence (human: 8.0%. LLM: 22.7%)
$\uparrow$ Universalism (human: 16.9%. LLM: 44.5%)
[…]
@@
Subsequent CHI-square tests with Bonferroni correction showed that LLM consistently shifted underlying values when rewriting people’s comments constructively across different cultures, topics, and stances; though effect sizes varied from small to medium (see Table 4.2). For instance, although Indians are known to prioritize Tradition (Sundberg et al., 1970), social Security and stability (Sharma, 2021), and Conformity to social norms (Anderson, 2012), LLM downplayed these values when rewriting their comments constructively. In contrast, in Western democratic contexts, such as the US where concern for others’ welfare (Universalism, Benevolence) is already high (Schwartz, 2007), LLM amplified these values even further.
On the other hand, in response to Islamophobic threads, many participants referred to Security concern around terrorism and extremism, which LLM deprioritized while highlighting Universalism and Benevolence during rewriting. Similarly, as observed in Phase 1 (see Figure 6B), opponents of same-sex marriage frequently appealed to Conforming to Traditional values, which the LLM minimized in its rewriting.
@@
According to Schwartz’s theory (Schwartz, 1994), the values that LLM consistently downplayed—Conformity, Security, and Tradition—fall under the higher-order value of Conservation (see Figure 3), which emphasizes stability and the preservation of existing social order. In contrast, Benevolence and Universalism—the values amplified by LLM—are part of the higher-order value Self-Transcendence, which prioritizes the welfare of others and the society.
@@
Notably, participants who wrote comments supporting same-sex marriage and Islam already emphasized Benevolence and Universalism by appealing to equality and rejecting dehumanization of these groups. This might explain why LLM did not shift underlying values or further amplified these values when rewriting comments from participants, who already held prosocial viewpoints (see Table 4.2). Moreover, our manual inspection revealed that for comments authored by opponents of same-sex marriage and Islam, changes in values in LLM-rewritten versions frequently (91% of cases) led to corresponding changes in stance—moving them toward more neutral or supportive positions. For instance, the example in Table 8 shows that although the participant expressed their views against same-sex marriage, LLM totally altered their arguments when rewriting constructively. In contrast, comments from those supportive did not undergo such changes.
@@
having gay marriage or legalizing it
@@
biologically there are only two gender
male and female. I don’t think gay
marriages should be legalized in India
as marriages are meant for having a
child which is not possible with gay
marriages and it will not only affect
the heritages but also the culture of India.
It’s important to understand the marriage is not exclusively
meant for procreation. Many heterosexual couples also choose
not to have children or can’t have them due to biological reasons.
However, they are still allowed to marry. Legalizing same-sex
marriages wouldn’t necessarily impact India’s culture or heritage,
as it merely extends the legal rights and protections of marriage
to all citizens. Moreover, gender diversity is biologically proven
beyond just male and female categories, adding another dimension
to our understanding of human nature.